OpenStudy (anonymous):

3/x - 3/x-1- 3/x-2? can someone please show me how to do this step by step please?

OpenStudy (anonymous):

its supposed to be rational expressions o.O

OpenStudy (anonymous):

-3(x+1)/x

OpenStudy (mathteacher1729):

1) Please be careful with parenthesis I assume you mean: 3/(x) - 3(x-1) - 3/(x-2) 2) Please be specific with your question... I assume you mean: " Express 3/(x) - 3(x-1) - 3/(x-2) as a single fraction. "

OpenStudy (mathteacher1729):

There are LOTS of examples here (scroll down to example 3 for problems exactly like this one) :) http://tutorial.math.lamar.edu/Classes/Alg/RationalExpressions.aspx

OpenStudy (anonymous):

there were no paranthesis in the problem :(

OpenStudy (mathteacher1729):

I'm guessing it was written as \[ \large \frac{3}{x}+\frac{3}{x-1}+\frac{3}{x-2}\] But when you write it in plain text you have to put whatever is in the denominator in a set of parenthesis.

OpenStudy (anonymous):

yes! and ohh okay